Three Velcro blocks, i.e Velcro on their ends, are
Answer #1

First Collision

We calculate the final velocity at the first collision:

P_0 = P_F conservation of linear momentum

m_1v_1 +m_2v_2 = (m_1+m_2)v

7.6*11+10*4 = (7.6+10)v

v = 7.0227 m/s

For the second collision:

Also apply the conservation of linear momentum:

P_0 = P_F

(m_1+m_2)v -m_3v_3= (m_1+m_2+m_3)V

V = 5.3518 m/s

For the spring we use the conservation of energy:

K = U

\frac{1}{2}(m_1+m_2+m_3)V^2 = \frac{kx^2}{2}

using given data:

x = 0.15137 m

C The lost energy is:

\Delta E = E_F-E_0 = \frac{kx^2}{2}-\left ( \frac{1}{2}m_1v_1^2+\frac{1}{2}m_2v_2^2+\frac{1}{2}m_3v_3^2 \right )

using given data:

\Delta E = -238.5 J